The width of a rectanglar plot is 4 feet longer than two-fifths the length. If the perimeter of the plot is 36 feet, find the length. How do I set up the equation?
Answer by algebrahouse.com(1659)

x = length
(2/5)x + 4 = width {width is 4 longer than two-fifths the length}

Perimeter of a rectangle is 2(width) + 2(length)

2x + 2[(2/5)x + 4] = 36 {substituted into perimeter formula}
2x + (4/5)x + 8 = 36 {used distributive property}
(14/5)x + 8 = 36 {combined like terms}
(14/5)x = 28 {subtracted 8 from each side}
x = 10 {divided each side by 14/5, which is multiplying by reciprocal, 5/14}

length is 10 feet

For more help from me, visit: www.algebrahouse.com